OpenStudy (paulalovee):

Pleasee help ! I will medal and fan I promise What is the approximate volume of a cylinder with a height of 2 ft and a radius of 6 ft? Use 3.14 to approximate pi. Express your answer in hundredths. ft3

OpenStudy (paulalovee):

@jabez177

OpenStudy (kendricklamar2014):

Formula for volume of a cylinder: \[V = \pi \times r^2 \times h\]

OpenStudy (kendricklamar2014):

Now just substitute and solve.

OpenStudy (paulalovee):

I got 226.08

jabez177 (jabez177):

And that is your answer. :)

OpenStudy (paulalovee):

Right or wrong ?

OpenStudy (paulalovee):

K ty ty ty

OpenStudy (kendricklamar2014):

Yup thats right

jabez177 (jabez177):

Anymore?